You long 1000 shares of ABC at $25 a share. Assume the Regulation T rate is...

  1. You long 1000 shares of ABC at $25 a share. Assume the Regulation T rate is 40%. You send the required amount to your broker. After a week ABC jumps to $27 in the market.

  1. Show your account.
  2. What is your buying power?
  3. Suppose you withdraw the maximum allowable cash from your account. Show your account.
Answer #1

Money needed to buy 1000 share of ABC @$25/share= 1000*25*40%=$10000

So, you have send $10000 to your broker.

a. Value of account= Initial Deposit+Profit/Loss=10000+(27-25)*1000=12000

b. buying power=12000/0.4=$30000

c. To hold the ABC share @$27/share you need =27*1000*40%=$10800

So, you can withdraw (12000-10800)=$1200 from your account and you will have 1000 share of ABC for $27/share